OpenStudy (anonymous):

Find the zeros of the function. f(x) = 4x3 - 12x2 - 40x

OpenStudy (anonymous):

You need to factor to find where the function goes to zero. Set f(x) = 0. Start by factoring the greatest common factor from f(x). Then try to rewrite your function as the product of binomials.

OpenStudy (cwrw238):

hint: greatest common factor is 4x

OpenStudy (anonymous):

4x(x^2-3x-10) what now?

OpenStudy (shinalcantara):

\[4x(x ^{2} - 3x - 10) = 0\] divide the whole equation by 4x to remove 4x

OpenStudy (aum):

\[ 4x(x^2-3x-10) = 4x(x-5)(x+2) = 0 \\ x = -2, ~~x =0, ~~x=5 ~~\text{are the roots.} \]

OpenStudy (anonymous):

thanks! @aum

OpenStudy (aum):

You are welcome.